OpenStudy (anonymous):

What is the point-slope form of the equation of a line that passes through the point (3, 5) and has a slope of 2? A. y + 3 = 2(x + 5) B y - 3 = 2(x - 5) C. y - 3 = 2(x + 5) D. y + 3 = 2(x - 5)

OpenStudy (anonymous):

A?

pooja195 (pooja195):

\[\huge~\rm~y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (anonymous):

M = 2 X1 = 5 y1 = 3

pooja195 (pooja195):

You flipped it (x,y)

OpenStudy (anonymous):

\[\huge~\rm~y-3_1=2(x-5_1)\]

OpenStudy (anonymous):

^ there

pooja195 (pooja195):

You can ignore the 1's lol \[\huge~\rm~y-3=2(x-5)\]

OpenStudy (anonymous):

Sooo its B?

pooja195 (pooja195):

Yep